OpenStudy (sloppycanada):

What is the difference between exact value and value? Example - Find the value of 11radians/6 I think it's .87

OpenStudy (sloppycanada):

@Michele_Laino

OpenStudy (michele_laino):

in order to find the measure of your angle in degrees, we have to solve this proportion: \[\pi :180 = \left( {11/6} \right):x\] where x is therequested value. If we apply the fundamental property of proportion, we get: \[x = \frac{{180 \times \frac{{11}}{6}}}{\pi } = ...\] please complete my computation

OpenStudy (sloppycanada):

x = 105.04

OpenStudy (michele_laino):

that's right!

OpenStudy (sloppycanada):

Okay so the value is the radians in degrees?

OpenStudy (michele_laino):

we have changed from radians to degrees, so we can say: 11/6 radians are equivalent to 105.04 degrees

OpenStudy (sloppycanada):

So it's asking me for the percentage. Got it. I guess the decimals would be approximate value.

OpenStudy (michele_laino):

yes! we can rewrite your answer in a different way, like this: \[\Large 105^\circ 2'4''\]

OpenStudy (michele_laino):

where as usual: 60' = 1 degree 60'' = 1'

OpenStudy (sloppycanada):

So the value of 11radians/6 is 105.04degrees. Assuming I round, 100 degrees?